# Generative Artificial Intelligence for Beamforming in Low-Altitude Economy

### v2

@@ Abstract
The growth of low-altitude economy (LAE) has driven a rising demand for efficient and secure communication. However, conventional beamforming optimization techniques struggle in the complex LAE environments. In this context, generative artificial intelligence (GenAI) methods provide a promising solution. In this article, we first introduce the core concepts of LAE and the roles of beamforming in advanced communication technologies for LAE. We then examine their interrelation, followed by an analysis of the limitations of conventional beamforming methods. Next, we provide an overview of how GenAI methods enhance the process of beamforming, with a focus on its applications in LAE. Furthermore, we present a case study using a generative diffusion model (GDM)-based algorithm to enhance the performance of aerial collaborative beamforming-enabled remote secure communications in LAE and simulation results demonstrate that our approach achieves 23% improvement in terms of secrecy rate and 18% reduction in energy consumption compared to baseline algorithms. Finally, promising research opportunities are identified.
@@ I Introduction
Low-altitude economy (LAE) is an economic model that relies on the utilization of low-altitude airspace for various commercial activities, with widespread applications in the fields such as precision agriculture, urban logistics, and disaster rescue [1]. By enabling efficient aerial resource allocation, LAE can quickly adapt to the demands of high-density users and dynamic tasks. Moreover, LAE enhances the service response speed and coverage by enabling rapid resource deployment, especially in remote or congested areas where traditional infrastructure may be limited or unavailable. However, to meet these demands, LAE faces challenging requirements on communication systems, to support efficient and flexible resource management in rapidly changing environments [2].
As a core technology in wireless communication systems, beamforming has been widely applied in LAE networking. By controlling the phase and amplitude of antenna arrays, beamforming enhances signal directionality and strength, thereby addressing the challenges such as signal coverage, interference suppression, and spectrum allocation. However, conventional beamforming optimization methods face limitations in LAE scenarios, which involve dynamic channel conditions, high-density interference, and complex task requirements. These methods struggle to adapt to real-time environmental changes and lack sufficient flexibility for multi-user collaboration and resource allocation. Therefore, relying solely on conventional beamforming optimization methods is insufficient to fully meet the communication demands of LAE networking.
@@
In contrast, generative AI (GenAI) models can inherently address these challenges. Specifically, GenAI models use the self-supervised learning enables diverse environment adaptation by capturing stochastic propagation phenomena and transferring knowledge across varied conditions, providing robust performance without environment-specific tuning. Furthermore, GenAI can synthesize realistic channel realizations and adapting to new situations with minimal training data. Additionally, uncertainty quantification mechanisms in generative models support optimal decision-making under limited channel state information, effectively balancing performance with data constraints in challenging LAE communication scenarios.
@@
By applying generative models, GenAI methods can capture complex channel characteristics and generate highly accurate channel state information (CSI), thereby improving the design and adjustment strategies for beamforming. Moreover, GenAI models present potential for adapting to frequent channel variations in dynamic environments and enhancing resource utilization and interference management. Crucially, GenAI models can learn from diverse data distributions, which enables them to evolve given various communication requirements and lays the foundation for multi-scenario applications within LAE.
@@
  • We propose a framework based on GenAI to optimize the performance of aerial collaborative beamforming-enabled remote secure communications in LAE. Simulation results show that the proposed framework significantly improves the secrecy rate and energy consumption of the considered system.
@@ II Overview of Low-altitude Economy Beamforming
In this section, we first introduce the core concepts of LAE. Next, we outline the fundamental principles and processes of beamforming. Finally, we explore the various applications of beamforming in LAE and emphasize its roles in communications and sensing.
@@ II-A1 Heterogeneous Communication Demands
The diverse operational requirements of LAPs in LAE scenarios create heterogeneous communication demands, thus bringing significant challenges in communication resource allocation. Specifically, in disaster reconnaissance and urban surveillance scenarios, LAPs must support high data rates to enable real-time video transmission [3]. In comparison, applications such as urban air taxis rely on ultra-reliable links with extremely low latency, as both reliability and responsiveness are critical for ensuring passenger safety. Meanwhile, for agricultural protection, LAPs prioritize wide coverage over transmission rate. Consequently, these varied and often conflicting requirements necessitate advanced communication solutions capable of dynamic resource allocation.
@@ II-A2 Dynamic Network Topologies
The mission-driven mobility of LAPs creates rapidly shifting network structures that challenge conventional communication paradigms. Specifically, large-scale agricultural spraying operations need real-time topology adaptation as LAP swarms cover vast farmlands [4]. Moreover, emergency response, such as wildfire monitoring, requires LAPs to dynamically form temporary relay chains for extended coverage in disaster zones. In addition, for urban air logistics, delivery LAP fleets execute sub-second topology switches for collision avoidance in congested corridors. Consequently, mission scenarios involving highly dynamic LAP mobility demand self-organizing networks capable of seamless topology transitions under strict latency constraints.
@@ II-A3 Complex Communication Environments
LAP operations face severe wireless signal disruption. Specifically, urban package delivery encounters unpredictable obstructions and reflections from buildings and alleys. Moreover, in precision crop dusting over terraced fields, agricultural LAPs suffer signal degradation and physical obstructions from dense foliage. Furthermore, event monitoring over crowded areas experiences intense interference from numerous nearby LAPs, phones, and other wireless devices. As a result, these environmental and interference challenges necessitate robust communication strategies adaptable to rapidly changing signal conditions for reliable and stable system performance.
Given the aforementioned characteristics and challenges of LAE, effective wireless signal management and optimization in low-altitude operations are crucial. Under this background, beamforming emerges as a relevant solution, which has substantial potential in tackling signal-related challenges.
@@ II-B Beamforming
Beamforming is a signal processing technique for optimizing the directionality of electromagnetic waves by controlling the phase and amplitude of each antenna element in an antenna array. To illustrate the key technologies involved in beamforming, Fig. 1 presents a timeline of representative methods. The details of these technologies are discussed below.
@@ II-B1 Beam Management
Beam management focuses on acquiring and maintaining reliable beam pairs to ensure robust directional communication.
@@
  • Beam Measurement: Beam measurement evaluates signal quality metrics including strength and interference across candidate beams.
@@ II-C2 Beamforming for Dynamic Network Topologies
Beamforming effectively addresses the challenge of constantly changing network connections in LAE by enabling rapid redirection of communication signals. For example, in scenarios such as multi-LAP emergency response operations, beamforming allows communication nodes to quickly shift their focused signal beams as LAPs relocate or reconfigure formations. Similarly, for large-scale LAP-based delivery systems, beamforming enables central communication hubs to dynamically track and maintain robust signal links with other LAPs during critical maneuvers such as takeoff, landing, and obstacle navigation. As a result, by continuously adjusting the beam direction towards moving targets, beamforming significantly improves connection stability and minimizes disruptions in these highly dynamic environments.
@@ II-C3 Beamforming for Complex Communication Environments
Beamforming provides a powerful solution for overcoming signal degradation in challenging low-altitude settings by concentrating signal energy along desired paths. Specifically, for LAPs operating in urban areas, beamforming helps overcome signal blockage caused by buildings by focusing the transmission towards the intended receiver, thereby finding stronger paths through or around obstacles, which improves link reliability. Moreover, in agricultural applications, beamforming effectively counters the signal weakening and physical obstruction caused by dense crops by directing a strong and focused beam, thus ensuring stable control and data transfer over longer distances. Furthermore, in crowded event areas with many LAPs and devices, beamforming minimizes interference by precisely directing signals towards intended receivers while reducing energy sent towards potential sources of interference. Consequently, this focused approach is essential for maintaining clear communication in congested and cluttered environments.
@@ II-D Applications of Beamforming in Low-Altitude Economy
Figure 2: Beamforming enables efficient communications and sensing applications in the LAE, such as interference avoidance, spectrum reuse, target tracking, and environmental monitoring.
@@ II-D2 Sensing
In addition to target tracking, another critical challenge is ensuring robust environmental sensing, particularly when wireless signals are prone to significant attenuation over long distances or when attempting to penetrate obstructions like debris or foliage. Moreover, such attenuation can severely degrade the quality of sensing data. To address this issue, beamforming mitigates attenuation effects by concentrating wireless signal energy toward areas of interest, thereby improving signal strength and increasing the fidelity of received data  [6]. For example, in LAE disaster response missions, concentrated beam scanning enables rapid detection of survivors hidden beneath rubble or obstructive terrain, thus enabling responders to efficiently locate individuals even in challenging environments where weak signals would render detection impossible.
@@ II-E3 Multi-Objective Optimization
Beamforming balances the interference reduction, energy efficiency, and communication reliability. However, beamforming often faces stringent environmental and resource constraints. Consequently, overlapping paths and limited spectrum in LAE scenarios make emergency communications and similar high-demand situations increasingly challenging. Therefore, addressing these multi-objective challenges calls for more advanced optimization techniques.
@@ II-E4 Antenna Technology Complexity
III Generative AI for Beamforming in Low-Altitude Economy
@@ III-A1 Variational Autoencoders (VAEs)
VAEs establish a probabilistic generative framework by explicitly optimizing a lower bound on data likelihood to learn latent representations that capture underlying data structures. Their distinctive advantage lies in naturally quantifying uncertainty through explicit variance modeling, while simultaneously ensuring consistent training convergence, despite typically generating reconstructions with less sharp details. As a result, the inherent uncertainty quantification makes VAEs particularly valuable for channel estimation in rapidly changing LAE conditions, where accurate beam prediction requires modeling the distribution of possible channel states rather than providing single point estimates. In light of these properties, the authors in [7] employed a VAE-based model to learn the complex probability distribution of CSI, thus parameterizing estimators that approach minimum mean square error performance even without perfect CSI during training phases. Moreover, by encoding channel characteristics into a latent space with explicit variance parameters, VAEs provide not only the channel estimates themselves but also confidence measures for those estimates, proving crucial when making communication decisions in variable channel conditions. In addition, the robustness of VAEs stems from this probabilistic generative structure. Specifically, by modeling a distribution over latent variables, VAEs inherently account for uncertainty and variability, thus promoting strong generalization and robust inference from partial, noisy, or degraded input, which is beneficial when sensor data is missing or corrupted.
@@ III-A2 Generative Diffusion Models (GDMs)
GDMs learn complex data distributions through a distinctive iterative denoising process. Specifically, GDMs simulate a gradual forward diffusion process that incrementally adds noise to data. Subsequently, the models learn to precisely reverse this process by progressively removing the noise step by step, thereby reconstructing the original data. Therefore, the step-wise denoising mechanism enables GDMs to capture the full transformation from pure noise to structured data, along with underlying statistical structures, thus leading to the generation of exceptionally high-fidelity and diverse samples. Moreover, this remarkable ability to generate accurate and varied data makes GDMs particularly valuable for modeling complex LAE channels, especially when detailed real-world channel measurements are limited or challenging to obtain. In light of these properties, in [8], the authors applied a GDM-based model to generate realistic MIMO channel samples conditioned on user location, even with limited real-world measurements. In addition, the robustness of GDMs is rooted in this iterative denoising procedure, thereby allowing them to correct arbitrary levels of corruption and achieve resilient reconstruction under strong jamming, complex noise, or signal dropout conditions, all common in low-altitude environments [9].
TABLE I: Generative AI Models for Beamforming Applications in Low-Altitude Economy
Model Channel Modeling & Estimation Resource Allocation Communication Security Energy Consumption Real-time Performance
@@
GDM Generate realistic MIMO channel samples via iterative denoising process [8] - Resist eavesdropping interference and mitigate location errors in beamforming [10] Iteratively refine beam patterns for energy-efficient power distribution Text: Fair Image: Fair Video: Fair
GAN Improve channel modeling and estimation [11], [12] - Suppress complex interference patterns through adversarial denoising [13] Simulate low-power beam patterns Text: Excellent Image: Excellent Video: Good
Transformer Identify relationships across modalities and enable parallel tracking for estimation Capture long-range dependencies and context for resource allocation Enhance secrecy via beamforming and spatiotemporal modeling [10] Optimize energy consumption by capturing environmental dependencies Text: Good Image: Good Video: Moderate
@@ III-A3 Generative Adversarial Networks (GANs)
GANs learn data distributions through an adversarial training framework, where a generator and a discriminator are trained in competition to implicitly capture underlying data structures. Crucially, the key strength of GANs lies in their ability to learn complex data patterns without needing explicit mathematical modeling of the data distribution. Therefore, the capability is particularly advantageous for interference suppression in LAE, which is characterized by non-stationary and complex interference patterns. Moreover, by training on clean signal data, GANs can learn what constitutes a clean signal and then use this knowledge to effectively remove interference from corrupted signals. As a result, GAN-based architectures enables robust interference mitigation even when the interference characteristics are unknown or constantly changing. For example, in [13], the authors proposed a GAN-based model for wireless signal denoising where the generator reconstructed clean signals from noisy inputs and the discriminator ensured that the outputs were indistinguishable from real clean signals, significantly improving clarity under complex noise and interference. In addition, the robustness of GANs arises from these adversarial training dynamics. Specifically, continuous competition forces the generator to approximate data distributions, including rare or extreme variants, which promotes strong resilience to distributional shifts and adversarial perturbations, frequently occurring in dynamic wireless environments.

### v3

@@ Abstract
The growth of low-altitude economy (LAE) has driven a rising demand for efficient and secure communications. However, conventional beamforming optimization techniques struggle in the complex LAE environments. In this context, generative artificial intelligence (GenAI) methods provide a promising solution. In this article, we first introduce the core concepts of LAE and the roles of beamforming in advanced communication technologies for LAE. We then examine their interrelation, followed by an analysis of the limitations of conventional beamforming methods. Next, we provide an overview of how GenAI methods enhance the process of beamforming with a focus on its applications in LAE. Furthermore, we present a case study using a generative diffusion model (GDM)-based algorithm to enhance the performance of collaborative beamforming-enabled secure communications in LAE and simulation results demonstrate that our approach achieves 23% improvement in terms of secrecy rate and 18% reduction in energy consumption compared to baseline algorithms. Finally, promising research opportunities are identified.
@@ I Introduction
Low-altitude economy (LAE) is an economic model that relies on the utilization of low-altitude airspace for various commercial activities, with widespread applications in fields such as precision agriculture, urban logistics and disaster rescue [1]. By enabling efficient aerial resource allocation, LAE can quickly adapt to the demands of high-density users and dynamic tasks. Moreover, LAE enhances the service response speed and coverage by enabling rapid resource deployment, especially in remote or congested areas where traditional infrastructure may be limited or unavailable. However, to meet these demands, the communication system of LAE faces challenging requirements for supporting efficient and flexible resource management in rapidly changing environments [2].
As a core technology in wireless communication systems, beamforming has been widely applied in LAE networking. By controlling the phase and amplitude of antenna arrays, beamforming enhances signal directionality and strength, thereby addressing the challenges such as signal coverage, interference suppression and spectrum allocation. However, conventional beamforming optimization methods face limitations in LAE scenarios, which involve dynamic channel conditions, high-density interference and complex task requirements. These methods struggle to adapt to real-time environmental changes and lack sufficient flexibility for multi-user collaboration and resource allocation. Therefore, relying solely on conventional beamforming optimization methods is insufficient to fully meet the communication demands of LAE networking.
@@
In contrast, generative AI (GenAI) models can inherently address these challenges. Specifically, GenAI models use self-supervised learning that enables adaptation to diverse environments by capturing stochastic propagation phenomena and transferring knowledge under varied conditions, thus providing robust performance without environment-specific tuning. Furthermore, GenAI can synthesize realistic channel realizations and adapt to new situations with minimal training data. Additionally, uncertainty quantification mechanisms in generative models support optimal decision-making under limited channel state information, effectively balancing performance with data constraints in LAE communication scenarios.
@@
By applying generative models, GenAI methods can capture complex channel characteristics and generate highly accurate channel state information (CSI), thereby improving the design and adjustment strategies for beamforming. Moreover, GenAI models present potential for adapting to frequent channel variations in dynamic environments and enhancing resource utilization and interference management. Crucially, GenAI models can learn from diverse data distributions, which enables them to evolve given various communication requirements within LAE.
@@
  • We propose a framework based on GenAI to optimize the performance of collaborative beamforming-enabled secure communications in LAE. Simulation results show that the proposed framework significantly improves the secrecy rate and energy consumption of the system.
@@ II Overview of Low-altitude Economy Beamforming
In this section, we first introduce the core concepts of LAE. Next, we outline the fundamental processes of beamforming. Finally, we explore the various applications of beamforming in LAE and emphasize its roles in communications and sensing.
@@ II-A1 Heterogeneous Communication Demands
The diverse operational requirements of LAPs in LAE scenarios create heterogeneous communication demands, thus bringing significant challenges in communication resource allocation. Specifically, in disaster reconnaissance and urban surveillance scenarios, LAPs must support high data rates to enable real-time video transmission [3]. In comparison, applications such as urban air taxis rely on ultra-reliable links with extremely low latency since both reliability and responsiveness are critical for ensuring passenger safety. Meanwhile, for agricultural protection, LAPs prioritize wide coverage over transmission rate. Consequently, these varied and often conflicting requirements necessitate advanced communication solutions capable of dynamic resource allocation.
@@ II-A2 Dynamic Network Topologies
The mission-driven mobility of LAPs creates rapidly shifting network structures that challenge conventional communication paradigms. Specifically, large-scale agricultural spraying operations need real-time topology adaptation as LAP swarms cover vast farmlands [4]. Moreover, emergency response like wildfire monitoring requires LAPs to dynamically form temporary relay chains for extended coverage in disaster zones. In addition, for urban air logistics, delivery LAP fleets execute sub-second topology switches for collision avoidance in congested corridors. Consequently, mission scenarios involving highly dynamic LAP mobility demand self-organizing networks capable of seamless topology transitions under strict latency constraints.
@@ II-A3 Complex Communication Environments
LAP operations face severe wireless signal disruption. Specifically, urban package delivery encounters unpredictable obstructions and reflections from buildings and alleys. Moreover, in precision crop dusting over terraced fields, agricultural LAPs suffer signal degradation and physical obstructions from dense foliage. Furthermore, event monitoring over crowded areas experiences intense interference from numerous nearby LAPs, phones and other wireless devices. As a result, these environmental and interference challenges necessitate robust communication strategies adaptable to rapidly changing signal conditions for reliable and stable system performance.
Given the aforementioned characteristics and challenges of LAE, effective wireless signal management and optimization in low-altitude operations are crucial. In this context, beamforming emerges as a relevant solution, which has substantial potential to tackle signal-related challenges.
@@ II-B Beamforming
Beamforming is a signal processing technique for optimizing the directionality of electromagnetic waves by controlling the phase and amplitude of each antenna element in an antenna array. As shown in Fig. 1, the beamforming-related technology details are discussed below.
@@ II-B1 Beam Management
Beam management focuses on acquiring and maintaining reliable beam pairs to ensure robust directional communications.
@@
  • Beam Measurement: Beam measurement evaluates signal strength and interference across all candidate beams.
@@ II-C2 Beamforming for Dynamic Network Topologies
Beamforming effectively addresses the challenge of constantly changing network connections in LAE by enabling rapid redirection of communication signals. For example, in scenarios such as multi-LAP emergency response operations, beamforming allows communication nodes to quickly shift their focused signal beams as LAPs relocate or reconfigure formations. Similarly, for large-scale LAP-based delivery systems, beamforming enables central communication hubs to dynamically track and maintain robust signal links with other LAPs during critical maneuvers such as takeoff, landing and obstacle navigation. As a result, by continuously adjusting the beam direction towards moving targets, beamforming significantly improves connection stability and minimizes disruptions in these highly dynamic environments.
@@ II-C3 Beamforming for Complex Communication Environments
Beamforming provides a powerful solution for overcoming signal degradation in challenging low-altitude settings by concentrating signal energy along desired paths. Specifically, for LAPs operating in urban areas, beamforming helps overcome signal blockage caused by buildings by focusing the transmission towards the intended receiver, thereby finding stronger paths through or around obstacles, which improves link reliability. Moreover, in agricultural applications, beamforming effectively counters the signal weakening and physical obstruction caused by dense crops by directing a focused beam, thus ensuring stable control and data transfer over longer distances. Furthermore, in crowded event areas with many LAPs and devices, beamforming minimizes interference by precisely directing signals towards intended receivers. Consequently, this focused approach is essential for maintaining clear communication in congested and cluttered environments.
@@ II-D Applications of Beamforming in Low-Altitude Economy
Figure 2: Beamforming enables efficient communications and sensing applications in the LAE, such as interference avoidance, spectrum reuse, target tracking and environmental monitoring.
@@ II-D2 Sensing
In addition to target tracking, another critical challenge is ensuring robust environmental sensing, particularly when wireless signals are prone to significant attenuation over long distances or when attempting to penetrate obstructions like debris or foliage. Moreover, such attenuation can severely degrade the quality of sensing data. To address this issue, beamforming mitigates attenuation effects by concentrating wireless signal energy toward areas of interest, thereby improving signal strength and increasing the fidelity of received data [6]. For example, in LAE disaster response missions, concentrated beam scanning enables rapid detection of survivors hidden beneath rubble or obstructive terrain, thus enabling responders to efficiently locate individuals.
@@ II-E3 Multi-Objective Optimization
Beamforming balances interference reduction, energy efficiency, and communication reliability. However, beamforming often faces stringent environmental and resource constraints. Consequently, overlapping paths and limited spectrum in LAE scenarios make emergency communications and similar high-demand situations increasingly challenging. Therefore, addressing these multi-objective challenges calls for more advanced optimization techniques.
@@ II-E4 Antenna Technology Complexity
II-E5 Channel Estimation Errors
In LAE beamforming, channel estimation errors lead to imprecise beam alignment, which degrades signal quality and increases the risk of signal leakage to eavesdroppers. To compensate for this, energy-constrained LAPs must increase transmission power, thus reducing the system lifetime. Moreover, this challenge is exacerbated as the number of collaborative entities increases since individual errors can accumulate and be magnified throughout the network.
@@ II-E5 Channel Estimation Errors
III GenAI for Beamforming in Low-Altitude Economy
@@ III-A1 Variational Autoencoders (VAEs)
VAEs establish a probabilistic generative framework by explicitly optimizing a lower bound on data likelihood to learn latent representations that capture underlying data structures. Their distinctive advantage lies in naturally quantifying uncertainty through explicit variance modeling, while simultaneously ensuring consistent training convergence, despite typically generating reconstructions with less sharp details. As a result, the inherent uncertainty quantification makes VAEs particularly valuable for channel estimation in rapidly changing LAE conditions, where accurate beam prediction requires modeling the distribution of possible channel states rather than providing single point estimates. In light of these properties, Baur et al. [7] employed a VAE-based model to learn the complex probability distribution of CSI, thus parameterizing estimators that approach minimum mean square error performance even without perfect CSI during training phases. Moreover, by encoding channel characteristics into a latent space with explicit variance parameters, VAEs provide not only the channel estimates themselves but also confidence measures for those estimates, proving crucial when making communication decisions in variable channel conditions. In addition, the robustness of VAEs stems from this probabilistic generative structure. Specifically, by modeling a distribution over latent variables, VAEs inherently account for uncertainty and variability, thus promoting strong generalization and robust inference from partial, noisy, or degraded input, which is beneficial when sensor data is missing or corrupted.
@@ III-A2 Generative Diffusion Models (GDMs)
GDMs learn complex data distributions through a distinctive iterative denoising process. Specifically, GDMs simulate a gradual forward diffusion process that incrementally adds noise to data. Subsequently, the models learn to precisely reverse this process by progressively removing the noise step by step, thereby reconstructing the original data. Therefore, the step-wise denoising mechanism enables GDMs to capture the full transformation from pure noise to structured data, along with underlying statistical structures, thus leading to the generation of exceptionally high-fidelity and diverse samples. Moreover, this remarkable ability to generate accurate and varied data makes GDMs particularly valuable for modeling complex LAE channels, especially when detailed real-world channel measurements are limited or challenging to obtain. In light of these properties, Wu et al. [8] applied a GDM-based model to generate realistic MIMO channel samples conditioned on user location, even with limited real-world measurements. In addition, the robustness of GDMs is rooted in this iterative denoising procedure, thereby allowing them to correct arbitrary levels of corruption and achieve resilient reconstruction under strong jamming, complex noise, or signal dropout conditions, all common in low-altitude environments [9].
TABLE I: GenAI Models for Beamforming Applications in Low-Altitude Economy
Model Channel Estimation Resource Allocation Communication Security Energy Consumption Real-time
@@
GDM Generate realistic channel samples [8] and refine noisy observations via iterative denoising [10] - Resist eavesdropping interference and mitigate location errors in beamforming [11] Iteratively refine beam patterns for energy-efficient power distribution Text: Fair Image: Fair Video: Fair
GAN Generate realistic samples to fill information gaps - Suppress complex interference patterns through adversarial denoising [12] Simulate low-power beam patterns Text: Excellent Image: Excellent Video: Good
Transformer Use a sequence of previous channel states to predict future channel parameter [13] Capture long-range dependencies and context for resource allocation Enhance secrecy via beamforming and spatiotemporal modeling [11] Optimize energy consumption by capturing environmental dependencies Text: Good Image: Good Video: Moderate
@@ III-A3 Generative Adversarial Networks (GANs)
GANs learn data distributions through an adversarial training framework, where a generator and a discriminator are trained in competition to implicitly capture underlying data structures. Crucially, the key strength of GANs lies in their ability to learn complex data patterns without needing explicit mathematical modeling of the data distribution. Therefore, the capability is particularly advantageous for interference suppression in LAE, which is characterized by non-stationary and complex interference patterns. Moreover, by training on clean signal data, GANs can learn what constitutes a clean signal and then use this knowledge to effectively remove interference from corrupted signals. As a result, GAN-based architectures enable robust interference mitigation even when the interference characteristics are unknown or constantly changing. For example, in [12], the authors proposed a GAN-based model for wireless signal denoising where the generator reconstructed clean signals from noisy inputs and the discriminator ensured that the outputs were indistinguishable from real clean signals, significantly improving clarity under complex noise and interference. In addition, the robustness of GANs arises from these adversarial training dynamics. Specifically, continuous competition forces the generator to approximate data distributions, including rare or extreme variants, which promotes strong resilience to distributional shifts and adversarial perturbations, frequently occurring in dynamic wireless environments.